Solve the Heat Equation du/dt=40d²u/dx² Subject to the conditions

u(0,t)=0, u(π,t)=0

u(x,0)=5sinx-2sin5x


1
Expert's answer
2021-02-04T14:18:18-0500

Solution:

We know that the solution of general heat equation

"\begin{array}{l}
\frac{\partial u}{\partial t}=k \frac{\partial^{2} u}{\partial x^{2}} \\
u(x, 0)=f(x) \quad u(0, t)=0 \quad u(L, t)=0
\end{array}"

is as follows:

"u_{n}(x, t)=B_{n} \sin \left(\frac{n \pi x}{L}\right) \mathbf{e}^{-k\left(\frac{n \pi}{L}\right)^{2} t} \quad n=1,2,3, \ldots" ...(i)

Now, given :

du/dt=40d²u/dx²

u(0,t)=0, u(π,t)=0

u(x,0)=5sinx-2sin5x

On comparing, we have:

"k=40,\ L=\pi,\ f(x)=5\sin x-2\sin 5x"

Choosing "n=1,\ B_n" as 5 and -2 respectively.

Putting these values in (i), we get,

"u(x, t)=2 \sin \left(\frac{ \pi x}{\pi}\right) \mathbf{e}^{-40\left(\frac{ \pi}{\pi}\right)^{2} t} -5 \sin \left(\frac{ \pi x}{\pi}\right) \mathbf{e}^{-40\left(\frac{ \pi}{\pi}\right)^{2} t}"

"\Rightarrow u(x, t)=2 \sin x \mathbf{e}^{-40 t} -5 \sin x \mathbf{e}^{-40 t}"